### v7

# Can LLM Agents Simulate Multi-Turn Human Behavior? Evidence from Real Online Customer Behavior Data

Abstract
Recent research shows that LLM Agents can generate “believable” human behaviors via prompt-only methods, and such agents have been increasingly adopted in downstream applications. However, existing evaluation of these agents only focuses on qualitative believability (whether human raters think they are accurate), leaving open questions of whether LLM agents can accurately generate step-by-step actions mimicking a particular human’s behavior in a multi-turn interaction task. In this work, we take shopping as a case study and present the first large-scale quantitative evaluation of state-of-the-art LLMs’ ability to accurately simulate human behavior. Using real-world data from 31,865 online shopping sessions containing 230,965 user actions, our evaluation reveals that prompt-based LLMs (DeepSeek-R1, Llama, Claude) achieve only 11.86% accuracy in generating human actions, highlighting a substantial gap in actual behavioral accuracy. Through experiments, we also showcase that strategies as simple as fine-tuning LLMs on real human click-through data augmented with synthesized reasoning traces can greatly enhance models’ performance. The fine-tuned Qwen2.5-7B achieves 17.26% action generation accuracy and 33.86% F1 score on final purchase prediction, representing substantial improvements of 5.4% and 13.85% over prompt-only baselines. This work establishes the first rigorous benchmark for human behavior simulation and provides actionable insights for developing more accurate LLM agents for future downstream applications.
ode]html
Can LLM Agents Simulate Multi-Turn Human Behavior? Evidence from Real Online Customer Behavior Data
Yuxuan Lu1,2, Jing Huang1, Yan Han1, Bingsheng Yao2, Sisong Bei1, Jiri Gesi1, Yaochen Xie1, Zheshen (Jessie) Wang1, Qi He1, Dakuo Wang1,2, 1Amazon.com, Inc., 2Northeastern University Correspondence: lu.yuxuan@northeastern.edu, d.wang@northeastern.edu
1 Introduction
Recent advances in large language models (LLMs) have enabled the simulation of human behavior across a range of applications, including web automation Gur et al. (2023); Zhou et al. (2024), social interaction behaviors Park et al. (2023), interpersonal trust behaviors Xie et al. (2024), and user interface interactions Taeb et al. (2024). These developments have sparked a growing interest in developing LLM Agent, agentic AI systems that uses LLMs to simulate human behaviors Chen et al. (2025a), in many domains, such as social science simulation Park et al. (2024), UI/UX testing Lu et al. (2025a), and expert annotation simulation Chen et al. (2025b). Specifically, in the online shopping domain, researchers have used LLM Agents as virtual customers to test website features Lu et al. (2025a), conduct automated A/B testing on design variants Wang et al. (2025), and evaluate agentic AI systems Sun et al. (2025). These applications build on the finding that LLM Agents are able to simulate “believable” human behavior Park et al. (2023, 2024), whether those behaviors are aligned with step-by-step human actions remains unclear. Thus, a fundamental question remains unanswered: How accurate can LLMs truly replicate human behavior?
In this paper, we focus specifically on the human behavior simulation task that is to generate the next action the user is most likely to perform in a multi-turn interaction session, based on the current observation and the history of past actions. For instance, in an online shopping scenario, the model observes the current webpage context (e.g., a product list) and the user’s action history (e.g., previous clicks or queries), and generates the next plausible action a human would take (e.g., add some product to the shopping cart).
Existing evaluations Park et al. (2023) of human behavior simulation primarily emphasize subjective measures of “believability” (“how much people feel it is like a human”) rather than the objective “accuracy” (“how much it acts like a human”). The most relevant works that measure the objective model accuracy focus only on the final outcome of a task (e.g., purchasing the final product or not Yao et al. (2023), or ultimately trusting the partner or not Xie et al. (2024)), without examining whether the intermediate decision and action sequences align with those of actual humans. Consequently, the field currently lacks a robust and quantitative understanding for assessing LLMs at the process-centric, action-level simulation of human behaviors.
To bridge this gap, we take online shopping as a case study and provide the first systematic evaluation of SOTA LLMs’ accuracy in process-centric, action-level behavior simulation tasks. We leverage a large-scale, real-world dataset consisting of 31,865 user click-through sessions from 3,526 users on an online shopping platform111Platform name omitted for anonymity. Each shopping session (Figure 1) comprises a series of timestamp-aligned context, actionpairs, where the context reflects the webpage observed by the user (e.g., product views, filter states), and the action denotes user inputs such as clicks, searches, or session termination actions. In total, the dataset has 230,965 user actions, and the final outcomes of the sessions include 4,432 purchase actions and 27,433 session termination actions. This dataset enables us to rigorously evaluate how accurately various LLMs can generate human-like behaviors at the action level. We evaluate different models on the next action generation task, benchmarking both the accuracy of generated actions throughout the session and the F1 score for final session outcome prediction (i.e. purchase or not), following protocols similar to existing work on modeling online shopping behavior.
Beyond evaluation, our dataset uniquely positions us to fine-tune LLMs to enhance their accuracy in behavior simulation tasks. While prior work has primarily relied on prompt-based approaches, we show that simply fine-tuning the model with user click-through data leads to significantly better accuracy in action generation and session outcome prediction. Furthermore, drawing inspiration from reasoning-augmented modeling DeepSeek-AI et al. (2025), we hypothesize that exposing models to intermediate reasoning traces–even if synthetically generated–can enhance their ability to simulate human behavior. To test this hypothesis, we augment our dataset with synthesized reasonings from action traces using Claude 3.5 Sonnet and fine-tune models using this augmented data (context, action, reasoning triplets) to learn how to generate not only accurate actions but also the underlying reasoning. Our results show that this reasoning-augmented fine-tuning further boosts model performance, highlighting the importance of modeling not just what humans do, but also why they do it. Taken together, our results provide the first quantitative evidence that out-of-the-box LLMs cannot accurately simulate action-level human behavior in realistic settings, and our ablation and error analyses provide insights to support future research.
In summary, this paper has three main contributions:
We conduct the first quantitative and process-centric evaluation of LLMs’ ability to simulate human web action behaviors using real-world online shopping data.
We show that out-of-the-box LLMs cannot accurately predict human behavior in the next action prediction setting, while simple fine-tuning achieves substantially better performance.
We demonstrate that fine-tuning LLMs with synthesized reasoning traces further enhances their accuracy, highlighting the importance of modeling not only what humans do but also why they do it for faithful human behavior simulation.
2 Related Works
2.1 Simulation of Human Behavior with LLM
The core function of the emerging LLM agent systems is their capability of generating human behaviors, in which a model takes a static user persona (e.g., preferences, demographics, or shopping habits), and the session data (e.g., a sequence of actions) as input to generate the next user action. Such systems have been extensively utilized and tested to simulate human behavior in a variety of scenarios. Park et al. (2023) simulated social behavior using generative agents in a virtual town, producing “believable” interactions. Xie et al. (2024) studied LLM agents in Trust Games to assess their ability to model human trust behavior. Park et al. (2024) used LLMs to simulate responses from 1,052 individuals in a social science survey. To simulate UI interaction, Lu et al. (2025a) proposed UXAgent, enabling LLMs to operate within web environments for simulated usability testing. Collectively, these studies underscore the growing potential of LLM-driven simulations to model and simulate complex, interpretable human behaviors.
However, the evaluation of these works remains limited in scope. Some focus on the subjective believability of process-centric action traces. For instance, Lu et al. (2025a) conducted qualitative interviews to assess participants’ perceptions of the realism of their UXAgent system. Similarly, Park et al. (2023) proposed an evaluation framework that identified emergent social behaviors among generative agents. On the other hand, works that pursue objective evaluation often do so in a single-shot, outcome-centric manner. Zhou et al. (2024) introduced WebArena, a controlled environment for benchmarking web agents based on task completion rates. ReAct Yao et al. (2023) measured success rates in simulation environments, overlooking the accuracy of model mimicing step-by-step human behavior. To date, no prior work has focused on objectively evaluating model-generated step-by-step actions–that is, assessing whether a model’s sequence of decisions faithfully aligns with human behavior at each step.
2.2 Reasoning in Human Behavior Simulation
Building on the chain-of-thought prompting strategy Wei et al. (2023), numerous studies have incorporated reasoning mechanisms into human behavior simulation. Park et al. (2023) pioneered agents equipped with reflection modules that synthesize memory and social context to support introspective decision-making. ReAct Yao et al. (2023) prompted models to generate reasoning traces and actions separately, improving task success rates in online shopping and gaming environments. Gur et al. (2023) proposed WebAgent, which uses a dedicated reasoning model to plan sub-steps in web browsing tasks, enhancing control and planning in real-world browser simulations. Beyond single-agent reasoning, systems such as ChatDev Qian et al. (2024) and RepoAgent Luo et al. (2024) adopt multi-agent setups, where agents with specialized roles (e.g., programmers, testers) engage in collaborative dialogues via structured prompts. These communicative exchanges support more robust collective reasoning, demonstrating how coordination between agents can improve the quality of generated reasoning traces.
However, the aforementioned works incorporate reasoning using prompt-only approaches for the action generation task. Whether reasoning can improve performance in fine-tuning settings remains an open question. Although datasets for reasoning behind model actions exist (e.g., for conversational agents (Dongre et al., 2025)), there is currently no ready-to-use dataset specifically designed for human behavior simulation that includes both human actions and the corresponding reasoning behind those actions. To address a similar cold-start problem in reinforcement learning, DeepSeek-AI et al. (2025) constructed a small-scale dataset of long reasoning traces by synthesizing examples through few-shot prompting, followed by reflection and verification. Inspired by their methodology, we apply a similar strategy to synthesize reasoning traces in the online shopping domain. This enables us to investigate whether integrating reasoning into fine-tuning can enhance a model’s ability to accurately simulate human behavior.
3 Method
3.1 Task Definition
In this section, we formally define the proposed human behavior simulation task: in the online shopping scenario, a shopping session is represented as a sequence of user actions , always starting with a search action and concluding with either a product purchase action or a termination action (i.e., the user closing the browser window).
At each time step , the model is tasked with generating both the reasoning and the action . The model input includes the current context (what the user currently observes), a sequence of previous contexts (what the user has observed) , a sequence of previous actions (what the user has done), and the corresponding synthetic reasoning trace (why the user had that action) within the same session. Formally, the model learns a function such that:
Observation Context
The Context (or “observation space”) of the web agent encompasses all available information on a webpage, including textual content, metadata, visual elements, and structural data. This context is designed to reflect how a human perceives and interprets a webpage, allowing the agent to process relevant features and perform tasks such as navigation, information retrieval, and interaction with page elements.
Previous research has explored various approaches to defining the context. Some methods rely on manually structured information parsing Yao et al. (2022), while others utilize raw HTML representations Gur et al. (2023) or accessibility tree embeddings Zhou et al. (2024). However, both approaches have notable limitations: manually structured information parsing requires significant human effort to develop parsing rules Yao et al. (2022), while raw HTML representations often contain extraneous information (e.g., CSS) that is not directly observable to human users, which may bias the LLM performance.
To ensure the adaptability to unseen websites, we define and implement a simplified HTML format as our context representation. This format removes non-relevant elements such as scripts, CSS, and purely visual components while preserving essential structural information. Using simplified HTML offers several advantages compared to custom defined formats (domain specific language, DSLs) or markdown-based representations: (1) important structural elements, such as lists and tables, remain intact, and (2) LLMs are already familiar with the HTML format, eliminating the need to redefine common elements like “button” and “input”.
[…]
Reasoning
[…]
Action
Previous research has explored various approaches to defining action spaces, including task-specific semantic actions such as “searching”, “adding items to a cart”, and “making purchases” Yao et al. (2022), as well as browser-level interactions like typing and clicking Lu et al. (2025a).
[…]
3.2 Synthesized Reasoning Trace
Reasoning traces are crucial for understanding users’ action choices but are difficult to collect; thus, they are often not available in behavioral datasets. We employ a reasoning synthesis pipeline to generate them using an LLM. To guide the reasoning generation process, we provide the LLM with the observation context and the corresponding action. Additionally, we record real human customers’ think-aloud shopping sessions Eccles and Arsal (2017) as in-context learning examples. We then prompt the LLM to generate a free-text reasoning explaining the user’s decision. Following recent works on reasoning-augmented LLMs Wei et al. (2023); DeepSeek-AI et al. (2025), the synthesized reasoning is not intended to replicate the actual human thought process. Instead, its purpose is to enhance the model’s predictive accuracy by providing structured intermediate representations that help the LLM better align actions with contextual cues. This approach ensures that the reasoning traces are coherent with the observed actions while improving the model’s behavioral fidelity and explainability.
3.3 Model Architecture
To incorporate these enriched action traces, we build on existing pre-trained LLMs as our base models. The input to the model consists of two components: (1) a sequence of historical contexts (what the user observed), and (2) the corresponding actions and generated reasoning trace (what the user did and why). The model is trained to generate the next reasoning corresponding action, enabling it to simulate human behavior with greater fidelity.
During the training stage, the model receives the full sequence of a user session—including context, synthetic reasoning, and action—as a single concatenated input. The training objective is to minimize the next-token prediction loss for the reasoning and action tokens, while the loss for the context tokens is masked out. Subsequently, in the evaluation stage, the model works in an autoregressive manner: it is provided with historical context and past reasoning traces and actions, and is asked to first generate the reasoning for the next action; then, based on the generated reasoning, it generates the next action.
4 Experiments
4.1 Dataset Construction
Our training dataset was constructed using data from one of the largest e-commerce platforms globally 222Platform name omitted for anonymity. The dataset contains 31,865 user sessions from 3,526 users in the online shopping scenario, comprising 230,965 user actions. The session’s final outcome includes 4,432 purchase actions and 27,433 session termination actions. We leveraged our data synthesis pipeline detailed in Section 3.2 to generate the synthetic reasoning for each action based on the context using Claude-3.5-Sonnet. The dataset was derived from traffic logs of a small group of users who explicitly opted into a beta testing feature and consented to data collection during the process. Users who opted out of data collection were excluded from the dataset. The data was anonymized to remove any personally identifiable information during processing. Due to the sensitive nature of the data and to protect user privacy, we are unable to release the dataset.
We extracted pairs of user actions and contextual information from the cleaned data. These raw logs were then structured into the standardized format defined in Section 3.1.
| Model | Generated Next Action | Session Outcome | ||||
| Accuracy | vs Base | v.s. DS-R1 | F1 Score | vs Base | v.s. DS-R1 | |
| Open-Source Models | ||||||
| DeepSeek-R1 | 11.86% | - | - | 20.01% | - | - |
| Llama 3.1 8B | 5.05% | - | -6.81% | 10.87% | - | -9.14% |
| Llama 3.1 70B | 8.19% | - | -3.67% | 12.69% | - | -7.32% |
| Mixtral 8x7B | 5.41% | - | -6.45% | 13.16% | - | -6.85% |
| Qwen2.5-7B | 4.25% | - | -7.61% | 11.94% | - | -8.07% |
| Qwen2.5-3B | 3.91% | - | -7.95% | 10.87% | - | -9.14% |
| Qwen2.5-1.5B | 3.27% | - | -8.59% | 7.94% | - | -12.07% |
| Mistral-7B-v0.3 | 4.25% | - | -7.61% | 11.27% | - | -8.74% |
| Llama 3.2 3B | 2.93% | - | -8.93% | 8.60% | - | -11.41% |
| Llama 3.2 1B | 3.71% | - | -8.15% | 3.09% | - | -16.92% |
| Proprietary Models | ||||||
| Claude 3.5 Haiku | 9.18% | - | -2.68% | 14.77% | - | -5.24% |
| Claude 3 Opus | 6.78% | - | -5.08% | 15.08% | - | -4.93% |
| Claude 3 Sonnet | 8.42% | - | -3.44% | 17.40% | - | -2.61% |
| Claude 3.5 Sonnet | 9.72% | - | -2.14% | 15.91% | - | -4.10% |
| Claude 3.5 Sonnet v2 | 11.69% | - | -0.17% | 18.54% | - | -1.47% |
| Claude 3.7 Sonnet | 9.34% | - | -2.52% | 12.81% | - | -7.20% |
| Fine-tuned Models | ||||||
| Qwen2.5-7B | 16.67% | - | 4.81% | 26.92% | - | 6.91% |
| + reasoning | 17.26% | 3.54% | 5.40% | 33.86% | 25.78% | 13.85% |
| Mistral-7B-v0.3 | 14.17% | - | 2.31% | 17.99% | - | -2.02% |
| + reasoning | 15.84% | 11.79% | 3.98% | 30.12% | 67.43% | 10.11% |
| Llama-3.2-3B | 9.31% | - | -2.55% | 4.73% | - | -15.28% |
| + reasoning | 15.77% | 69.39% | 3.91% | 33.99% | 618.60% | 13.98% |
4.2 Evaluation and Metrics
Evaluation Dataset
We used a subset of the dataset that was not used during training as the test set, ensuring that no user sessions in the test set were seen by the model during fine-tuning. To create test cases, we took the second and all subsequent actions within each session, excluding the first action, since it lacks any preceding context. For each test case, the model was provided with the historical context, along with all previous actions and corresponding synthetic reasoning traces in the same session. The model is first tasked with predicting the next reasoning, and then, based on the reasoning it generated, it produces the corresponding action.
Prompt-Only Baseline
To assess pre-trained models’ ability in predicting human behavior, we evaluated a set of instruction-tuned LLMs under the in-context learning (ICL) setting (a.k.a., prompt-based setting). Specifically, we used several variants of Claude, LLaMA, and Mistral as representatives for general-purpose pre-trained LLMs, along with DeepSeek-R1 as a representative for reasoning LLMs. To preserve user privacy, we run open-source model on our own GPU cluster and
In this setup, each model was provided with the historical context and previous user actions from the session, and prompted to generate both the next reasoning and the next action. The generated actions were used to compute macro accuracy for evaluation. These baselines reflect the commonly adopted approach of using powerful LLMs without domain-specific fine-tuning, allowing us to directly assess the impact of fine-tuning on realistic human behavior simulation.
Evaluation Metrics
We evaluate model performance across two key dimensions: Next Action Generation and Session Outcome Classification.
For Next Action Generation, we evaluate whether the model can accurately generate user actions during the process. We use an exact match accuracy, where a generation is considered correct only if the action type, action target (e.g., search box or product link), and action attribute (such as search keyword) exactly match the ground truth. To avoid skewing the results toward longer sessions, we compute per-session accuracy first and then average across sessions, ensuring each session contributes equally to the final score.
[…]
4.3 Experimental Setup
We fine-tuned multiple language models to evaluate their performance on the action generation task. The models used in our experiments include:
Fine-Tuned Models: Different versions of Llama 3.2, Qwen 2.5, and Mistral.
Baseline Models: Different versions of Claude, Llama, Mistral, and DeepSeep-R1.
All fine-tuned models were trained using the same dataset and pipeline to ensure a fair comparison.
Model training was performed on a GPU cluster consisting of NVIDIA H200 GPUs, with each training job utilizing eight nodes, each equipped with eight GPUs, for a total of 64 GPUs, each with 140 GB of GPU memory. A typical job takes 3 hours on 64 GPUS, and in total, we used about 3700 H200 GPU hours for our experiments.
We employed Fully Sharded Data Parallel Zhao et al. (2023) for efficient training. All sequences were padded or truncated to a context length of 40k tokens. We used a per-device batch size of 1, resulting in a global batch size of 64. The learning rate was set to 2e-5 with a cosine scheduler for adaptive learning rate adjustment. Models were trained for 1 epoch. For example, training Mistral-7B-v0.3 with a 40k token context window in our setup requires approximately 130GB of GPU memory per GPU, which is the largest model we can train.
4.4 Evaluation Results and Analysis
Following previous works Lutz et al. (2024); Deng et al. (2023), we evaluate the performance of prompt-based LLMs in simulating human behavior on shopping scenarios, with results presented in Table 1. Our findings indicate that while state-of-the-art LLMs have demonstrated strong capabilities in various tasks and domains, their ability to accurately simulate human behaviors remains limited, with Claude 3.5 sonnet v2 achieving 11.69% accuracy in next action prediction. Compared to general instruction-tuned models, reasoning-focused DeepSeek-R1 achieved a higher accuracy of 11.86% in the action generation task, suggesting that incorporating reasoning mechanisms offers some advantage. Similarly, for final outcome prediction, DeepSeek-R1 achieved an F1 score of 20.01%, outperforming other baseline models. Additionally, larger model usually performs better than their smaller variants. These results indicate that reasoning ability positively impacts performance on human behavior prediction tasks, supporting the hypothesis that such tasks benefit from models trained with reasoning-oriented objectives.
We then fine-tuned various smaller open-sourced LLMs, including LLaMA 3.2, Qwen 2.5, and Mistral, using our training dataset. The results demonstrate that fine-tuning LLMs with action traces and synthesized reasoning traces significantly enhances performance. Qwen 2.5-7B achieved 17.26% accuracy in action generation, significantly surpassing DeepSeek-R1 by 5.4% (, McNemar’s test). Similarly, LLaMA 3.2-3B reached an F1 score of 33.99% on the final outcome prediction task, further confirming the effectiveness of fine-tuning. Additionally, all fine-tuned models significantly outperformed their own ICL variants (, McNemar’s test). These findings underscore that incorporating domain-specific fine-tuning with synthesized reasoning traces leads to substantial improvements in the accuracy of human online shopping behavior simulation.
Figure 2 compares the distribution of actions generated by different models with real human behavior. Human users rarely apply filters and instead rely heavily on iterative search, averaging 2.82 searches per session, which is over seven times more frequent than filter actions. This reflects natural behaviors such as correcting typos and revising keywords (as shown in Sec. 4.6). In contrast, pre-trained LLMs such as Claude 3.5 Sonnet and DeepSeek-R1 tend to stick to the initial search keyword without revision, overuse filtering actions, and produce disproportionately high purchase rates that diverge from actual user behavior. We hypothesize that this bias arises because existing LLM agent benchmarks like WebShopYao et al. (2022) and WebArenaZhou et al. (2024) primarily evaluate task completion (i.e., making a purchase), which incentivizes models to optimize for purchase-heavy trajectories rather than realistic browsing and search behavior.
The fine-tuned models exhibit a distribution that more closely aligns with human action patterns, capturing a more natural balance between search refinement, product clicks, and minimal reliance on filtering. Unlike pre-trained models, they do not prematurely converge on purchase-oriented behavior and instead reflect the exploratory and iterative nature of real user sessions.
| Example 1 | Example 2 | |
| Previous Action | search for “disney gift” | search for “tee conector” |
| Human Next Action | search for “disney gift card” | search for “tee connector” |
| Qwen-2.5-7B | search for “disney gift card” | search for “tee connector” |
[…]
4.5 Ablation Study
To evaluate the impact of synthesized reasoning traces on the model’s action generation capability, we conducted an ablation experiment. In the base setting, we removed reasoning traces from both the training and evaluation stages to isolate the contribution of reasoning-augmented learning. This setup allows us to directly assess whether exposure to synthesized reasoning traces improves the model’s ability to generate user actions and predict outcomes.
From Table 1, most models exhibited substantial improvements in action generation accuracy when trained with reasoning traces, with relative gains ranging from 3.54% to 69.69%. Similar benefits were observed for final outcome prediction, where F1 scores increased by 25.78% to over 600% across models. For example, Qwen2.5-7B achieved a 33.86% F1 score with reasoning traces but dropped to 26.92% without them, underscoring the importance of explicit reasoning guidance. These results confirm that incorporating synthesized reasoning traces enhances both step-wise action generation and final outcome modeling, reinforcing their value in fine-tuning models for human behavior simulation.
4.6 Error Analysis
We analyze model behavior across different error types, comparing two ICL models, Claude, representing general-purpose chat models, and DeepSeek-R1, representing reasoning-augmented models, with a fine-tuned compact model, Qwen2.5–7B. We focus on five distinct error types. Didn’t terminate, didn’t click, and didn’t search indicate cases where the model chose a different action instead of terminating, clicking, or searching, as the human user did. Searched wrong keyword refers to instances where the model performed a search like the user, but used a different (and incorrect) search query. Clicked wrong button captures cases where the model clicked, but on a different button than the one selected by the human user. Illegal actions generated by models are excluded from this analysis. Overall, Claude and DeepSeek-R1 exhibit very similar error profiles, suggesting that the reinforcement learning process used to incorporate reasoning introduces similar biases to those seen in standard chat-based ICL models. A major shared failure mode of ICL models is their tendency to continue sessions or make the purchase even when human users would have terminated them by closing the browser window. This aligns with earlier findings that ICL agents are more likely to complete purchases, likely because current LLMs are optimized to fulfill task goals rather than follow subtle social cues or termination heuristics. In contrast, fientuned Qwen2.5–7B demonstrates more accurate alignment with human termination behavior.
We also observe that Qwen-2.5–7B better captures iterative search behaviors commonly exhibited by real users, such as retrying a query with corrected keywords or fixing typos after an unsatisfactory result. In contrast, ICL-based models like Claude tend to persist with the original query result and proceed to a different type of action, rather than issuing a refined search. As shown in Table 2, Qwen-2.5–7B more closely matches human revisions in both examples, whereas Claude chooses to visit the product shown on the current search result page. These findings suggest that fine-tuning not only reduces the overall error rate but also enhances the model’s ability to capture fine-grained user behavior patterns, such as correcting typos, retrying failed searches, and deciding when to terminate a session.
5 Discussion and Future Works
5.1 Action Misalign Between Human and LLM Agents
Existing research has shown that large language models (LLMs) can generate highly “believable” human behavior simulations, supporting various interactive and social simulation scenarios. However, our results indicate that general-purpose pretrained LLMs, despite their strong ability in a variety of tasks and applications, struggle to accurately generate user actions. This is reflected in their performance—DeepSeek-R1 and Claude 3.5 Sonnet v2, for example, achieved only 11.86% and 11.69% accuracy, respectively, on the next action generation task. These findings suggest that while LLMs may appear human-like and produce plausible and believable behaviors, accurate step-wise next action prediction requires additional fine-tuning and explicit alignment with real human actions.
Building on this observation, we find that current LLM agents are often misaligned with real human users, particularly in the online shopping domain. For example, as shown in Figure 2 and Figure 3, out-of-the-box LLMs tend to make more purchases than human users. We hypothesize that this stems from the training objectives of many LLMs, which are typically optimized for the final task completion (e.g., sending emails or making purchases) and evaluated based on task completion and task efficiency (i.e., completing the task with a minimal number of steps). To enable accurate human behavior simulation, it is essential to close this gap and ensure models capture what a real user will do under certain circumstances.
We also advocate for the development of evaluation metrics that reflect the irrational nature of human behavior. On one hand, many user actions are rational and goal-directed, where a single correct next step exists and can be evaluated against a gold standard. On the other hand, users also exhibit irrational or weakly deterministic behaviors, such as semi-random clicks, and typographical errors that even the same person may not replicate consistently. A robust metric should not penalize plausible variants that remain consistent with human trajectories and should explicitly assess an agent’s ability to reproduce human-like error patterns. For example, the capacity to generate human-like typos is also essential for applications like automatic UI/UX evaluation Lu et al. (2025b), where assessing typo correction features requires agents to reproduce the kinds of errors real users make. A human-aligned evaluation should therefore distinguish true errors from acceptable variability and report performance across both rational and non-rational action regimes.
5.2 Reasoning in Next Action Generation
Our experiments further highlight the critical role of synthesized reasoning traces in improving action generation. Removing reasoning traces from the training data led to a notable performance drop across all models. Additionally, most models showed measurable improvements when trained with reasoning traces, reinforcing the importance of providing explicit reasoning signals. These results suggest that reasoning traces not only enhance model interpretability but also serve as a guiding mechanism, enabling the model to make more contextually appropriate and human-aligned decisions.
Overall, our findings emphasize the necessity of domain-specific fine-tuning using enriched, context-aware human behavioral data. By incorporating synthesized reasoning traces, we bridge the gap between the believable yet imprecise predictions of general LLMs and the accurate behavioral simulations required for real-world applications.
5.3 Future Works
Future research should move beyond treating reasoning traces as static supervision and instead focus on actively enhancing the model’s reasoning abilities. One direction is to use reinforcement learning to improve the reasoning for action prediction accuracy, where better reasoning is defined by its ability to produce more accurate action prediction. Another direction is to train models to generate reasoning that more closely reflects human cognitive processes, so the reasoning itself can serve as a source of qualitative insight when analyzing or interacting with LLM agents. In addition to advancing reasoning, future work could scale up the dataset to improve robustness, incorporate user personas for more personalized simulations, and incorporate user intents with intent detection methods. Finally, incorporating vision-language models (VLMs) may enhance the model’s understanding of graphical interfaces in complex web environments.
6 Conclusion
In this work, we present the first quantitative, process-centric evaluation of state-of-the-art large language models (LLMs) for simulating human behavior in an online shopping task. Our study demonstrates that LLMs fine-tuned with real-world human behavioral data and synthesized reasoning traces showed a significant enhancement in their ability to generate user actions. Compared to in-context learning and pretrained baselines, our method achieved substantial improvements, including a 13.98% absolute gain in F1 score for final outcome prediction and a 5.4% increase in action generation accuracy.
These results underscore the critical role of explicit reasoning traces in aligning model predictions with human behavior. By enriching behavioral datasets with structured reasoning, we move closer to accurate and interpretable simulations of human behavior. Our findings highlight a promising direction for the development of LLM agent systems capable of producing realistic and explainable human-like behaviors in online shopping and other interactive domains.
Limitations
[…]
Appendix A Prompts
Reasoning Synthesize Prompt:
Baseline model evaluation prompt: